# Imitation Learning-Based Online Time-Optimal Control with Multiple-Waypoint Constraints for Quadrotors

### v1

@@ Abstract
Over the past decade, there has been a remarkable surge in utilizing quadrotors for various purposes due to their simple structure and aggressive maneuverability, such as search and rescue, delivery and autonomous drone racing, etc. One of the key challenges preventing quadrotors from being widely used in these scenarios is online waypoint-constrained time-optimal trajectory generation and control technique. This letter proposes an imitation learning-based online solution to efficiently navigate the quadrotor through multiple waypoints with time-optimal performance. The neural networks (WN&CNets) are trained to learn the control law from the dataset generated by the time-consuming CPC algorithm and then deployed to generate the optimal control commands online to guide the quadrotors. To address the challenge of limited training data and the hover maneuver at the final waypoint, we propose a transition phase strategy that utilizes polynomials to help the quadrotor ’jump over’ the stop-and-go maneuver when switching waypoints. Our method is demonstrated in both simulation and real-world experiments, achieving a maximum speed of $7$ $m/s$ while navigating through $7$ waypoints in a confined space of $6.0$ $m\times 4.0$ $m\times 2.0$ $m$ [video [^3] https://youtube.com/playlist?list=PLrUtAo-zaR9xmOzuae2mXLVsNZxzyAgWa&si=1WNQfKzKF6gG36dK]. The results show that with a slight loss in optimality, the WN&CNets significantly reduce the processing time and enable online optimal control for multiple-waypoint-constrained flight tasks.
@@ I INTRODUCTION
Quadrotors are one of the most agile unmanned aerial vehicles, which have immense potential for time-critical tasks such as delivery, inspection, search and rescue operations, etc [1]. However, due to challenges in navigation, trajectory generation, and control, few quadrotors are able to execute aggressive flight in real-world applications. Over the past decade, many researchers have been dedicated to addressing these challenges, aiming to push the boundaries of quadrotor applications.
@@
In this letter, we discuss the scenario where the quadrotor has to autonomously fly through multiple waypoints (including dynamic waypoints whose trajectories are unknown in advance) aggressively with time-optimal trajectories. We present an imitation learning-based approach and develop a neural network called WN&CNets (waypoint-constrained navigation and control network) to generate control commands online given the waypoints and the states of the quadrotor. Our approach can make quadrotors replan in real-time with high frequency to handle multiple waypoints including dynamic waypoints and also the deviation caused by disturbance or model inaccuracy.
Figure 1: A quadrotor navigates through $7$ waypoints online by the proposed WN&CNets in real-world experiments, which achieves a maximum speed of $7$ $m/s$ in a space of $6.0$ $m\times 4.0$ $m\times 2.0$ $m$.
@@
Common approaches for aggressive trajectory generation can generally be classified into three categories: polynomial-based methods, sampling-based methods and optimization-based methods. Polynomial-based methods [2, 3] leverage the quadrotors’ differential flatness property [4] to efficiently calculate trajectories subject to waypoint constraints. However, due to the inherent smoothness of polynomials, quadrotors are unable to reach their input limits consistently, which leads to suboptimal performance [5]. Sampling-based methods [6, 7, 8] search in the state spaces until available trajectories are found or specific trajectories are considered and the best one is selected. However, the high dimensionality of quadrotors’ state space poses a challenge when conducting these searches within the limited time. Simplifying the quadrotor’s dynamic model is a common approach to increase the solving efficiency but it will lose part of the quadrotors’ maneuverability and the time optimality. Therefore, optimization-based methods may emerge as a viable approach to avoid sacrificing the optimality. State-of-the-art optimization algorithms such as the CPC [5], are capable of calculating optimal trajectories for a quadrotor navigating through multiple waypoints and can be extended to multi-drone scenarios [9]. However, these optimization algorithms typically rely on numerical methods, which can be time-consuming. Thus, they are impractical for situations where online calculation and replanning are needed, such as aggressive flight in dynamic environments, mission changing during the flight or autonomous drone racing, etc.
To achieve online aggressive flight, deep reinforcement learning (RL) [10, 11] is utilized to learn time-optimal control policy for navigating specific racing tracks. Model Predictive Contouring Control (MPCC) [12, 13] is also developed to track pre-defined time-optimal trajectories with disturbances, tracking errors and changing waypoints, etc. However, these methods either rely on time-consuming offline computation or are only valid in very specific environments.
@@
Imitation learning presents an alternative online approach by using neural networks to learn pre-generated trajectory databases, allowing the quadrotors to fly towards the goal from different states [14, 15, 16, 17]. However, these works do not consider multiple waypoints or are limited to two-dimension spaces. Recently, the Guidance & Control Networks [18] have been trained for tasks involving two upcoming waypoints, but the waypoints’ position is very limited and there are relatively large errors when flying through the waypoint.
@@
In this letter, we propose a novel neural network, the WN&CNets, to learn the control law in optimal trajectories with $2$ waypoints, which are generated by the CPC method [5] (Fig. 2). Based on the original WN&CNets, we also develop a transition phase strategy to help the quadrotor ’jump over’ the stop-and-go maneuver introduced by the WN&CNets due to the limited computing resources to let the quadrotor consecutively fly multiple waypoints aggressively without hovering.
@@
    propose a polynomial-based transition phase strategy to help the quadrotor ’jump over’ the stop-and-go maneuver when switching waypoints while ensuring that the quadrotor can keep near time-optimal performance with very limited training data.
@@
    validate and analyze our approach both in simulation and real-world flight tests where the quadrotor navigates through $7$ waypoints with a maximum speed of $7$ $m/s$ in a space of $6.0$ $m\times 4.0$ $m\times 2.0$ $m$ to demonstrate the feasibility of the proposed method.
@@ II-A Quadrotor dynamics
In this letter, we utilize the same quadrotors’ dynamics model as the one previously presented in the work of [5].
${\dot{\mathbf{x}}}=\mathbf{f}_{dyn}({\mathbf{x}},{\mathbf{u}})=\begin{cases}{\mathbf{v}}\\ \mathbf{g}+\frac{1}{{m}}\mathbf{R}({\mathbf{q}}){{\mathbf{u}}}\\ \frac{1}{2}\Lambda({\mathbf{q}})\begin{bmatrix}0\\ {\boldsymbol{\omega}}\end{bmatrix}\\ {\mathbf{J}}^{-1}({\boldsymbol{\tau}}-{\boldsymbol{\omega}}\times{\mathbf{J}}{\boldsymbol{\omega}})\end{cases}$ (2)
where
${{\mathbf{u}}}=\begin{bmatrix}0\\ 0\\ \sum{{u}_{i}}\end{bmatrix}$
is the thrust vector of the quadrotor and
[…]
is the torque vector of the quadrotor. In the equations above ${\mathbf{v}}$, ${\mathbf{q}}$ and ${\boldsymbol{\omega}}$ represent velocity, quaternions and angular velocity, respectively. $\mathbf{R}({\mathbf{q}})$ is the rotation matrix. $\mathbf{J}$ is the inertia matrix. ${l}$ is the arm length. ${u_{i}}$ are the thrust of the rotors, which are the inputs of the dynamic system (2).
@@ II-B The optimization problem
The CPC problem for a specific planning task is formulated on a set of discrete time segments, with the number of the time segments $N$ determined based on the estimation of the flight task distance. At node $k$, denote the system states as $\mathbf{x}_{k}$, where $k\in[0,N]$. The system states include dynamic states and progress states, which can be presented as
@@
Because minimizing flight time is the only goal of the optimization problem, the problem can be described as
@@
The constraints of the optimization include
  1. 1.
    dynamic constraints
    $\begin{cases}\mathbf{x}_{k+1}-\mathbf{x}_{k}=dt\cdot f_{dyn}(\mathbf{x}_{k},\mathbf{u}_{k})\\ dt=t_{N}/N\end{cases}$ (4)
  2. 2.
    boundary constraints
    $\begin{cases}\mathbf{x}_{0}=\mathbf{x}_{init}\\ \mathbf{x}_{N}=\mathbf{x}_{end}\end{cases}$ (5)
  3. 3.
    input constraints
    $\begin{cases}\mathbf{u}_{min}-\mathbf{u}_{k}\leq 0\\ \mathbf{u}_{k}-\mathbf{u}_{max}\leq 0\end{cases}$ (6)
  4. 4.
    subsequence constraints
    $\begin{cases}{\mathbf{\lambda}_{k+1}}-{\mathbf{\lambda}_{k}}+{\mu_{k}}=0\\ \mathbf{\lambda}_{0}-1=0,\mathbf{\lambda}_{N}=0\\ -\mu_{k}\leq 0\\ \mathbf{\lambda}_{k}^{j}-{\mathbf{\lambda}_{k}^{j+1}}\leq 0,\forall k\in[0,N),j\in[0,M)\\ \end{cases}$ (7)
    where M is the number of waypoints.
  5. 5.
    complementary progress constraints
@@
    where $d_{tol}$ is the defined upper bound of the tolerance range, and ${\mathbf{P}}_{wj}$ is the position of the $j^{th}$ waypoint.
@@ II-C Dataset generation
The computational complexity of solving the optimization problem above makes it impossible to run the solver online during the flight. To tackle this problem, we use the WN&CNets to map the quadrotor’s waypoints, current states and the next step states to form an online model predictive controller (MPC). However, to generate the dataset, traversing all the states and the waypoints becomes impossible with the increasing quadrotor’s state space and the number of waypoints. Thus, with limited computational resources, to generate the dataset for the WN&CNets, in this letter, we only generate the time-optimal trajectories from hover states to hover states while passing through $2$ waypoint (the latter one is a hover state waypoint), which has $N+1$ nodes and node $k$ $(k\in[0,N])$ contains the quadrotor’s state at $t_{k}$, including relative position to the waypoints $\Delta\mathbf{p}_{k}$, velocity $\mathbf{v}_{k}$, acceleration $\mathbf{a}_{k}$, rate $\boldsymbol{\omega}_{k}$, and thrust $T_{k}$. Notably, the original CPC method [5] uses quaternions to represent attitudes and provides a general acceleration as the thrust command. However, the relationship between quaternions and the control commands exhibits a high degree of nonlinearity, which poses challenges during the training process. In this work, instead of quaternions and thrust, we use accelerations along the three axes ($\mathbf{a}_{k}=\frac{1}{{m}}\mathbf{R}({\mathbf{q}_{k}}){{\mathbf{u}_{k}}}$), as they are an equivalent representation.
To form an MPC, we can take $\Delta\mathbf{p}_{k}$, velocity $\mathbf{v}_{k}$, acceleration $\mathbf{a}_{k}$, rate $\boldsymbol{\omega}_{k}$ as the inputs of the network and $\boldsymbol{\omega}_{k+1}$ and thrust $T_{k+1}$ as the output of the neural network to form the state-control pairs, that is
$\begin{bmatrix}\Delta\mathbf{p}_{k}&\mathbf{v}_{k}&\mathbf{a}_{k}&\boldsymbol{\omega}_{k}\end{bmatrix}\rightarrow\begin{bmatrix}\boldsymbol{\omega}_{k+1}&T_{k+1}\end{bmatrix}$ (11)
@@
$\begin{bmatrix}\Delta\mathbf{p}_{k-d}&\mathbf{v}_{k-d}&\mathbf{a}_{k-d}&\boldsymbol{\omega}_{k-d}\end{bmatrix}\rightarrow\begin{bmatrix}\boldsymbol{\omega}_{k+1}&T_{k+1}\end{bmatrix}$ (14)
@@ II-D WN&CNet architecture and training
The inputs for the WN&CNets consist of the quadrotor’s relative positions to the waypoints ($\Delta\mathbf{p}\in\mathcal{R}^{3\times 2}$), velocity ($\mathbf{v}\in\mathcal{R}^{3}$), acceleration ($\mathbf{a}\in\mathcal{R}^{3}$), and angular rates ($\boldsymbol{\omega}\in\mathcal{R}^{3}$) of node $k-d$. The corresponding outputs of the WN&CNets are the angular rates ($\boldsymbol{\omega}_{c}\in\mathcal{R}^{3}$) and thrust commands ($T_{c}$) of node $k+1$ which will be sent to the quadrotors.
@@
The general network architecture employed in this work is depicted in Fig. 2. For the training process, we utilize the mean absolute error (MAE) as the loss function, as it demonstrates superior training effects compared to the mean squared error (MSE). The dataset is divided into training data (80%) and validation data (20%) to assess the performance of the model. We employ the Adam optimizer to update the weights of the networks. The accuracy achieved by the trained networks is up to $99.80\%$ after 40 epochs. The loss decreases quickly during the training progress and finally reaches about $0.04$.
II-E Polynomial-based transition phase strategy
@@
As stated above, to decrease the computational burden, in our dataset, we only have $2$ waypoints and the second one is a hover-state waypoint. So the quadrotor needs to hover at the second waypoint. Thus, for a multiple-waypoint flying task, a waypoint-switching strategy is needed to switch the waypoint during the flight. An intuitive waypoint-switching strategy is that when the quadrotor arrives at the second waypoint, it switches the next two waypoints as its target resulting in a stop-and-go maneuver, which violates the time-optimal purpose.
To solve this issue, we need a transition phase to a) help the quadrotor jump over the stop-and-go maneuver, b) keep near time-optimal performance, and c) run online. In this work, we choose polynomials to help the quadrotor ’jump over’ the stop-and-go maneuver.
@@
Figure 3: For same direction flights, the velocity is used to determine the entrance and exit of the transition phase. Polynomials are used to jump over the stop-and-go maneuver.
An important challenge lies in determining when to enter and exit the transition phase. We discuss our strategy in $2$ scenarios. Let $\mathbf{wp}_{i}$ ($i=1,2,3$) represent the position vector of the $i^{th}$ waypoint. The first scenario is shown in Fig. 3 where $(\mathbf{wp}_{2}-\mathbf{wp}_{1})(\mathbf{wp}_{3}-\mathbf{wp}_{2})>0$, which we call ’same direction flight’. After waypoint 1, with the WN&CNets, the quadrotor will decelerate to hover at waypoint 2 and then accelerate to waypoint 3. Its velocity curve resembles a ’v’ shape which leads to a stop-and-go maneuver. To avoid this maneuver, we set a parameter $v_{in}$ as a threshold to determine the entrance of the transition phase and we denote the corresponding point by $p_{in}$. Similarly, we set another threshold $v_{out}$ to determine the exit of the transition phase and the corresponding point is denoted by $p_{out}$. At the point where the predicted velocity of the quadrotor is larger than the threshold $v_{p}>v_{out}$, the quadrotor exits the transition phase. It should be noted that $v_{p}$ is the predicted velocity using the WN&CNets to steer the quadrotor from the hover state (waypoint 2) toward the following waypoints.
@@
Figure 4: For opposite direction flights, use acceleration to determine the entrance and exit of the transition phase.
@@
When the boundaries of the polynomials are determined (positions, velocities, and accelerations), the last key element to determine a polynomial is the traverse time. In this letter, we estimate the traverse time of the polynomial by the uniform acceleration assumption. We assume that within the transition phase, the quadrotor has a constant acceleration and then the corresponding traverse time can be calculated.
@@
Figure 5: With the proposed transition phase, the quadrotors don’t have to hover at the waypoints.

### v2

@@ Abstract
Over the past decade, there has been a remarkable surge in utilizing quadrotors for various purposes due to their simple structure and aggressive maneuverability, such as search and rescue, delivery and autonomous drone racing, etc. One of the key challenges preventing quadrotors from being widely used in these scenarios is online waypoint-constrained time-optimal trajectory generation and control technique. This letter proposes an imitation learning-based online solution to efficiently navigate the quadrotor through multiple waypoints with time-optimal performance. The neural networks (WN&CNets) are trained to learn the control law from the dataset generated by the time-consuming CPC algorithm and then deployed to generate the optimal control commands online to guide the quadrotors. To address the challenge of limited training data and the hover maneuver at the final waypoint, we propose a transition phase strategy that utilizes MINCO trajectories to help the quadrotor ’jump over’ the stop-and-go maneuver when switching waypoints. Our method is demonstrated in both simulation and real-world experiments, achieving a maximum speed of $5.6m/s$ while navigating through $7$ waypoints in a confined space of $5.5m\times 5.5m\times 2.0m$ [video [^3] https://youtube.com/playlist?list=PLrUtAo-zaR9xmOzuae2mXLVsNZxzyAgWa&si=1WNQfKzKF6gG36dK]. The results show that with a slight loss in optimality, the WN&CNets significantly reduce the processing time and enable online optimal control for multiple-waypoint constrained flight tasks.
@@ I INTRODUCTION
Quadrotors are one of the most agile unmanned aerial vehicles, which have immense potential for time-critical tasks such as delivery, inspection, search and rescue operations, etc [1]. Over the past decade, many researchers have been dedicated to addressing the challenges in navigation, trajectory generation and control, aiming to push the boundaries of quadrotors to execute aggressive flight in real-world applications [2, 3].
Figure 1: The quadrotor continuously passes through moving target points at high agility with the proposed WN&CNets and the MINCO transition.
@@
In this letter, we discuss the scenario where the quadrotor has to autonomously fly through multiple waypoints (including dynamic waypoints whose trajectories are unknown in advance) aggressively with time-optimal trajectories in a confined space, e.g. drone racing challenges [4] and indoor rescue [5]. We present an imitation learning-based approach and develop a neural network called WN&CNets (waypoint-constrained navigation and control network) to generate control commands online given the waypoints and the states of the quadrotor. Our approach can make quadrotors replan in real-time with high frequency to handle multiple waypoints including dynamic waypoints and also the deviation caused by disturbance or model inaccuracy.
@@
Common approaches for aggressive trajectory generation can generally be classified into three categories: polynomial-based methods, sampling-based methods and optimization-based methods. Polynomial-based methods [6, 7] leverage the quadrotors’ differential flatness property [8] to efficiently calculate trajectories subject to waypoint constraints. However, due to the inherent smoothness of polynomials, quadrotors are unable to reach their input limits consistently with polynomial trajectories at finite order or with finite segments of polynomials [1]. Sampling-based methods [9, 10] search in the state spaces until available trajectories are found or specific trajectories are considered and the best one is selected. However, the high dimensionality of quadrotors’ state space poses a challenge when conducting these searches within the limited time. Simplifying the quadrotor’s dynamic model is a common approach to increase the solving efficiency but it will lose part of the quadrotors’ maneuverability and the time optimality. Optimization-based methods emerge as a viable approach to avoid sacrificing optimality. State-of-the-art optimization algorithms such as the CPC [11], are capable of calculating optimal trajectories for a quadrotor navigating through multiple waypoints and can be extended to multi-drone scenarios [12]. However, these optimization algorithms typically rely on numerical methods, which can be time-consuming. Thus, they are impractical for situations where online calculation and replanning are needed, such as aggressive flight in dynamic environments, etc.
@@
To achieve online time-optimal flight with multiple-waypoint constraints, various innovative planning and control methods are developed and combined for autonomous drone systems [13, 14, 15]. Model Predictive Contouring Control (MPCC) [16, 17] is developed to track pre-defined trajectories while providing time-optimal online replanning. Deep reinforcement learning (RL) [18, 19] is utilized to train control policy for navigating racing tracks while coping with model uncertainty. Imitation learning also presents an alternative online approach by using neural networks to learn optimal trajectories [20, 21, 22]. However, these works do not consider multiple waypoints or are limited to two-dimension spaces. Recently, the Guidance & Control Networks [23] have been trained for tasks involving two upcoming waypoints, but the waypoints’ position is very limited and there are relatively large errors when flying through the waypoint.
@@
In this letter, we propose a novel neural network, the WN&CNets, to learn the control law in completely time-optimal trajectories with $2$ waypoints, which are generated by the CPC method [11] (Fig. 2). Based on the original WN&CNets, we also apply MINCO trajectories [13] as a lightweight tool to develop a transition phase strategy to help the quadrotor ’jump over’ the stop-and-go maneuver introduced by the WN&CNets, due to the limited computing resources. With the aid of the MINCO transition, the WN&CNets can empower the quadrotor to consecutively fly through multiple waypoints in a confined space, performing remarkably aggressive and agile flight maneuvers without hovering.
@@
    propose a MINCO-based transition phase strategy to help the quadrotor ’jump over’ the stop-and-go maneuver when switching waypoints while ensuring that the quadrotor can keep near time-optimal performance with very limited training data.
@@
    validate and analyze our approach both in simulation and real-world flight tests where the quadrotor navigates through $7$ waypoints with a maximum speed of $5.6m/s$ in a space of $5.5m\times 5.5m\times 2.0m$ and fly with complicated trajectories (up to 22 waypoints) to demonstrate the feasibility of the proposed method.
@@ II-A Quadrotor dynamics
In the equations above ${\mathbf{p}}$, ${\mathbf{v}}$, ${\mathbf{q}}$ and ${\boldsymbol{\omega}}$ represent position, velocity, attitude and angular velocity, respectively.
In this letter, we utilize the same quadrotors’ dynamics model as the one previously presented in the work of [11], which includes a linear drag term [8] represented by an approximated diagonal matrix $\mathbf{D}=diag(d_{x},d_{y},d_{z})$.
${\dot{\mathbf{x}}}=\mathbf{f}_{dyn}({\mathbf{x}},{\mathbf{u}})=\begin{cases}{\mathbf{v}}\\ \mathbf{g}+\frac{1}{{m}}\mathbf{R}({\mathbf{q}}){{\mathbf{u}}}-\mathbf{R}({\mathbf{q}})\mathbf{D}\mathbf{R}^{T}({\mathbf{q}})\mathbf{v}\\ \frac{1}{2}\Lambda({\mathbf{q}})\begin{bmatrix}0\\ {\boldsymbol{\omega}}\end{bmatrix}\\ {\mathbf{J}}^{-1}({\boldsymbol{\tau}}-{\boldsymbol{\omega}}\times{\mathbf{J}}{\boldsymbol{\omega}})\end{cases}$ (2)
Readers are referred to [11] for the details of the meaning of the symbols.
@@ II-B The optimization problem
The CPC problem for a specific planning task is formulated on a set of discrete time segments, with the number of the time segments $N$ determined based on the estimation of the flight task distance. The system states include dynamic states and progress states, which can be presented as
@@
As minimizing flight time is the only goal of the optimization problem, the problem can be described as
@@
The constraints of the optimization include dynamic constraints, boundary constraints, input constraints, subsequence constraints and complementary progress constraints, which are based on the original CPC method [11]. For the complementary progress constraints,
@@
where $d_{tol}$ is the defined upper bound of the tolerance range, and ${\mathbf{P}}_{wj}$ is the position of the $j^{th}$ waypoint. If we only consider the time optimality without heading constraints, we use the same constraints (4) as in [11]. However, for some flight tasks where the heading at the waypoint needs to be specified, we will revise constraint (4) to
$\begin{cases}{\mu_{k}^{j}}(\left\|{\mathbf{P}_{k}}-{\mathbf{P}}_{wj}\right\|_{2}^{2}+(\psi_{k}-\psi_{wj})^{2}-{\nu}_{k}^{j}):=0\\ -\nu^{j}_{k}\leq 0\\ \nu^{j}_{k}-d^{2}_{tol}\leq 0\end{cases}$ (5)
where $\psi_{k}$ is the heading at node $k$ and $\psi_{wj}$ is the desired heading at the $j^{th}$ waypoint. In this letter, we focus on generating trajectories which are as fast as possible, and thus do not impose any constraints on the heading. However, in the next section, some experimental attempts and analysis are conducted on using heading as one of the target states for waypoints.
@@ II-C Dataset generation
The computational complexity of solving the optimization problem above makes it impossible to run the solver online during the flight. To tackle this problem, we use the WN&CNets to map the quadrotor’s waypoints, current states and the next step states to form an online controller. However, to generate the dataset, traversing all the states and the waypoints becomes impossible with the increasing quadrotor’s state space and the number of waypoints. Thus, with limited computational resources, to generate the dataset for the WN&CNets, in this letter, we only generate the time-optimal trajectories from hover states to hover states while passing through $2$ waypoint (the latter one is a hover state waypoint). Each time-optimal trajectory has $N+1$ nodes and node $k(k\in[0,N])$ contains the quadrotor’s state at $t_{k}$, including relative position to the waypoints $\Delta\mathbf{p}_{k}$, velocity $\mathbf{v}_{k}$, attitude $\mathbf{q}_{k}$, rate $\boldsymbol{\omega}_{k}$, and thrust $T_{k}$. Notably, the original CPC method [11] uses quaternions to represent attitudes and provides a general acceleration as the thrust command. However, the relationship between quaternions and the control commands exhibits a high degree of nonlinearity, which poses challenges during the training process. Instead of quaternions, we use roll, pitch and yaw to represent the orientation in this work.
To be more specific, we can take $\Delta\mathbf{p}_{k}$, velocity $\mathbf{v}_{k}$, attitude $\mathbf{q}_{k}$ and thrust $T_{k}$ as the inputs of the network and $\boldsymbol{\omega}_{k+1}$ and thrust $T_{k+1}$ as the output of the neural network to form the state-control pairs, that is
$\begin{bmatrix}\Delta\mathbf{p}_{k}&\mathbf{v}_{k}&\mathbf{q}_{k}&T_{k}\end{bmatrix}\rightarrow\begin{bmatrix}\boldsymbol{\omega}_{k+1}&T_{k+1}\end{bmatrix}$ (6)
@@
$\begin{bmatrix}\Delta\mathbf{p}_{k-d}&\mathbf{v}_{k-d}&\mathbf{q}_{k-d}&T_{k-d}\end{bmatrix}\rightarrow\begin{bmatrix}\boldsymbol{\omega}_{k+1}&T_{k+1}\end{bmatrix}$ (7)
@@ II-D WN&CNet architecture and training
The inputs for the WN&CNets consist of the quadrotor’s relative positions to the waypoints ($\Delta\mathbf{p}\in\mathcal{R}^{3\times 2}$), velocity ($\mathbf{v}\in\mathcal{R}^{3}$), attitude ($\mathbf{q}\in\mathcal{R}^{3}$), and thrust ($T$) of node $k-d$. The corresponding outputs of the WN&CNets are the angular rates ($\boldsymbol{\omega}_{c}\in\mathcal{R}^{3}$) and thrust commands ($T_{c}$) of node $k+1$ which will be sent to the quadrotors.
@@
The general network architecture employed in this work is depicted in Fig. 2. For the training process, we utilize the mean squared error (MSE) as the loss function. We employ the Adam optimizer to update the weights of the networks. We build the network and conduct training using the PyTorch framework. The loss decreases quickly during the training progress and reaches about $0.04$ after 40 epochs.
II-E MINCO-based transition phase strategy
@@
As stated above, to decrease the computational burden, in our dataset, we only have $2$ waypoints and the second one is a hover-state waypoint. So the quadrotor needs to hover at the second waypoint. Thus, for a multiple-waypoint flight task, a waypoint-switching strategy is needed to switch the waypoint during the flight. An intuitive waypoint-switching strategy is that when the quadrotor arrives at the second waypoint, it switches the next two waypoints as its target resulting in a stop-and-go maneuver, which violates the time-optimal purpose.
To solve this issue, we need a transition phase to a) help the quadrotor jump over the stop-and-go maneuver, b) keep near time-optimal performance, and c) run online. In this work, we choose MINCO trajectories [13] to help the quadrotor ’jump over’ the stop-and-go maneuver. MINCO is a novel minimum control effort optimization framework with efficient computation. By parameterizing the quadrotor’s flat output trajectory, it can eliminate constraints on system dynamics, initial state, and final state while ensuring precise waypoint traversal. We apply it as a lightweight tool to compute polynomials while minimizing traversal time. We directly give the waypoints and don’t need to construct safe flight corridors (SFC) in the framework.
@@
(a) For same direction flights, the velocity is used to determine the entrance and exit of the transition phase. MINCO trajectories are used to jump over the stop-and-go maneuver.
@@
(b) For opposite direction flights, use acceleration to determine the entrance and exit of the transition phase.
Figure 3: Illustrative sketch of the CPC and its deficiency in multi-drone racing scenarios.
An important challenge lies in determining when to enter and exit the transition phase. We discuss our strategy in $2$ scenarios. Let $\mathbf{wp}_{i}$ ($i=1,2,3$) represent the position vector of the $i^{th}$ waypoint. The first scenario is shown in Fig. 3 (a) where $(\mathbf{wp}_{2}-\mathbf{wp}_{1})(\mathbf{wp}_{3}-\mathbf{wp}_{2})>0$, which we call ’same direction flight’. After waypoint 1, with the WN&CNets, the quadrotor will decelerate to hover at waypoint 2 and then accelerate to waypoint 3. Its velocity curve resembles a ’v’ shape which leads to a stop-and-go maneuver. To avoid this maneuver, we set a parameter $v_{in}$ as a threshold to determine the entrance of the transition phase and we denote the corresponding point by $p_{in}$. Similarly, we set another threshold $v_{out}$ to determine the exit of the transition phase and the corresponding point is denoted by $p_{out}$. At the point where the predicted velocity of the quadrotor is larger than the threshold $v_{p}>v_{out}$, the quadrotor exits the transition phase. It should be noted that $v_{p}$ is the predicted velocity using the WN&CNets to steer the quadrotor from the hover state (waypoint 2) toward the following waypoints.
@@
Figure 4: The sketch of the waypoint-switch strategy: with the proposed transition phase, the quadrotors don’t have to hover at the waypoints.